Find the demand functions for the individuals below, the budget constraint is pxx+pyy=m

23

Bill:U(x ,x )=x y 12

2/5 3/5 Buster: U(x, y)=x y

23 Ben: U(x, y) = (x+1) (y+2)

Barbara: U(x ,x ) =3x +2y 12

Beth: U(x, y)=min{x , y}


1
Expert's answer
2020-08-14T16:48:27-0400

Bill:


"U(x,y)=x^2y^3"

"m=12"

"MU_x=\frac { \delta U}{ \delta x}=2xy^3"

"MU_y=\frac { \delta U}{ \delta y}=3x^2y^2"

"\frac {MU_x}{p_x}=\frac {MU_y}{p_y}"

"xp_x=\frac {2}{3}yp_y"

"xp_x+yp_y=12"

"yp_y=9.2"

"xp_x=4.8"

Buster:


"U(x,y)=x^{\frac{2}{3}}y^{\frac {3}{5}}"

"m=23"

"MU_x=\frac { \delta U}{ \delta x}= (\frac {x}{y})^{\frac {2}{5}}"

"MU_y=\frac { \delta U}{ \delta y}=(\frac {y}{x})^{\frac {3}{5}}"

"\frac {MU_x}{p_x}=\frac {MU_y}{p_y}"


"yp_x=xp_y"